study the following sketch of a molecular orbital


Solving limiting reactant problems in solution Suppose 0.407 g of sodium acetate is dissolved in 100. mL of a 44.0 m Maqueous solution of ammonium sulfate. Calculate the final molarity of sodium cation in the solution. You can assume the volume of the it. solution doesnt change when the sodium acetate is dissolved n Be sure your answer has the correct number of significant digits ALL

Suppose 0.407 g of sodium acetate is dissolved in 100. mL of a 44.0 mM aqueous solution of ammonium sulfate. 
Answer #1

moles of sodium acetate (CH?COONa) = 0.407/82 = 0.00496 mole

so, moles of sodium cation = 0.00496 mole

final volume of solution = 100 ml = 0.1 L

so final molarity of sodium cation = 0.00496/0.1 = 4.96 x 10^-2 M

if you satisfied please like it.. thanks...